### v1

@@
Abstract
Recently, there has been an increased interest in NeRF methods which reconstruct differentiable representation of three-dimensional scenes. One of the main limitations of such methods is their inability to assess the confidence of the model in its predictions. In this paper, we propose a new neural network model for the formation of extended vector representations, called uSF, which allows the model to predict not only color and semantic label of each point, but also estimate the corresponding values of uncertainty. We show that with a small number of images available for training, a model quantifying uncertainty performs better than a model without such functionality. Code of the uSF approach is publicly available at https://github.com/sevashasla/usf/.
Keywords:
Neural semantic field Neural network Learning Uncertainty 3D-scene.
@@ 1 Introduction
Figure 1: We propose the model named uSF to predict both color and semantic labels and estimate corresponding uncertainty.
Over the past few years, many researchers have focused on the development of the field of differentiable scene representation. This development began with the appearance of the first NeRF method [9], which demonstrated the photo-realistic quality of novel views generation. Various modifications of the original method allow us to additionally solve such problems as depth estimation [18], semantic segmentation[20, 6], object detection [4]. The disadvantages of NeRF methods include high computational complexity, the need for a large number of images in the training dataset and the inability to estimate the confidence of the model in its predictions. Quantifying uncertainty is a crucial aspect for minimizing possible risks for tasks in various fields, for example robotics and autonomous driving.
As shown in Fig. 1 we propose a novel NeRF-based method that reconstructs the neural radiance and the neural semantic fields. The previously proposed NeRF methods only considered estimating the uncertainty for color predictions, i.e. regression task. Our method additionally introduces the possibility of quantifying uncertainty for predicted semantic labels, i.e. classification task. There are two types of uncertainty: aleatoric and epistemic. Aleatoric uncertainty reflects the noise inherent in the input data. Epistemic uncertainty reflects what the model does not know due to a lack of data, too simple architecture, or inefficient training process. In both cases, aleatoric uncertainty is estimated. The high value of uncertainty may also indicate errors occurred in the data preparation process, for example, poor quality of labelling in segmentation tasks. We choose aleatoric uncertainty over epistemic uncertainty because quantifying aleatoric uncertainty does not require major changes in network architecture and multiple inferences or training. Adding uncertainty estimation to the algorithm pipeline does not degrade the quality of reconstruction with a sufficient number of images and improves the quality with a small training dataset. Our method is also computationally efficient because it uses trainable positional encoding based on hashing [10]. We summarize our main contributions as follows:
  • We have proposed a new neural network model called uSF, which allowes to add the ability to quantify the uncertainty for predicted neural semantic field, while maintaining quantifying uncertainty for neural radiance field.
@@ 2.1 Neural semantic fields
There are NeRF methods that can predict not only the color at each point of the scene, but also corresponding semantic labels. The Semantic-NeRF [20] method reconstructs the neural semantic field with the help of an additional head in the model architecture. The process of getting semantic labels at a point is similar to getting a color. Such process is highly robust to the presence of outliers in the training data, yielding quality results even with 90% noise.
Another approach to segmentation is shown in the NeRF-SOS [2] method. The work successfully combines self-supervised 2D visual features for 3D semantic segmentation. Separate head outputs for segmentation are trained similar to Dino-ViT [1]. NeSF [17] is the method that produces generalizable neural semantic field. The authors propose to firstly recover 3D density fields upon which 3D semantic segmentation model is trained supervised by posed 2D semantic maps. The pre-trained NeRF model is used to sample the volumetric density grid for further conversion into a semantic-feature grid. Semantic Ray [6] also has generalization ability and uses Cross-Reprojection Attention module to obtain such result.
@@ 2.2 Uncertainty estimation
To estimate the aleatoric uncertainty, a probability distribution is imposed on the outputs of the model. This approach assumes that there is noise in the input data, the distribution parameters of which can be predicted as some function of the data. To predict epistemic uncertainty, it is necessary to impose a distribution on the model parameters (Bayesian neural networks (BNN) [7, 11, 3]). The disadvantage of these methods for quantifying epistemic uncertainty is the requirement to train the model several times and make multiple inferences. There are also approaches that simultaneously predict both types of uncertainties [evidential_regress, 13, 5].
Uncertainty quantification in various implementations has already been applied in several NeRF models. So in the Active NeRF  [12] method, the estimated uncertainty value is used when adding new frames in iterations of active learning. In the NeRF-W  [8] method, uncertainty is used to identify non-static objects that are removed from the prediction. In the S-NeRF  [15] method, it is assumed that all possible neural radiance fields for the scene are distributed according to a prior distribution, on the basis of which and the input data the model predicts a posterior distribution, with which the uncertainty is estimated. The CF-NeRF[14] method works similarly, but does not impose any preliminary restrictions on the distribution.
@@ 3.1 Neural Network Model
Figure 2: uSF architecture. Orange rectangles are Linear + ReLU layers. Yellow rectangles are outputs of our model. We study different variants of our architecture in terms of the color (c0-c3) and the semantic (s0-s1) branches. In these variants, only features taken for uncertainty predictions change.
The architecture of our model is MLP with ReLU activation functions and a hidden state dimension equal to 64. Similar to Active NeRF [12] pipeline we add an additional head to predict the color variance $\beta_{rgb}$ which we use as rgb uncertainty. In our work we assume that it depends on the view direction.
We add two heads for semantic segmentation task. One outputs mean values $(sem_{1},\dots,sem_{n})$, and the other predicts the variance $\beta_{sem}^{2}$. We consider outputs of semantic heads to depend only on coordinates of the point, but not on view direction. We have implemented these assumptions in our model.
@@ 3.2 Uncertainty for neural radiance field
Figure 3: From left to right we show the predicted color, semantic labels, rgb uncertainty and semantic uncertainty during training process
@@ 3.3 Uncertainty for neural semantic field
$\widehat{p_{i}}=\mathbb{E}SoftMax(\{logit_{1},\dots,logit_{n}\})_{i}$ (2)
@@
During calculation we use the <<reparametrization trick>>. Let $\xi\sim\mathcal{N}(a,\sigma^{2})$, and $\eta\sim\mathcal{N}(0,1)$. Then $\xi\leavevmode\nobreak\ {\buildrel d\over{=}}\leavevmode\nobreak\ a+\sigma\cdot\eta$. Then to generate logits, we firstly generate random values from $\mathcal{N}(0,1)$, and then convert them to the form we want by multiplying by the standard deviation ($\beta_{sem}$) and adding the mean ($sem_{i}$). The example of the predicted semantic uncertainty is shown in the Fig. 3
@@ 3.4 Learning approach
$L_{uncert}=\frac{1}{N}\sum_{i=1}^{N}\left(\frac{||C(r_{i})-rgb_{i}||^{2}}{2\beta_{rgb}(r_{i})^{2}}+\frac{1}{2}\log{\beta_{rgb}(r_{i})^{2}}+\frac{\eta}{N_{si}}\sum_{j=1}^{N_{si}}\alpha(r_{i}(t_{j}))\right).$ (7)
@@
$lr(i)=lr(0)\cdot 0.1^{\min(1,\frac{i}{N_{i}})}\cdot\min\left(1,10^{-3}+\frac{i}{N_{wi}}\right),$ (8)
@@
We use an early stopping during training. We firstly calculate mean of last 3 LPIPS values on eval dataset. Then if |current $|\text{current LPIPS}-mean(\text{last 3 LPIPS})|<10^{-3}$ is two times in a row, the training process is stopped.
@@ 4.1 Hardware and software setup
We have performed experiments on NVIDIA GeForce 2080 Ti, NVIDIA GeForce 2080, NVIDIA GeForce 3060 graphics cards. The model takes about 6GB of video memory during training. According to our estimates, full training takes approximately 4-5 hours for one scene.
@@ 4.2 Datasets and environments
We use a pre-processed dataset from the Semantic NeRF [20] method with image size of $640\times 480$. Each scene contains 900 images with different view directions. We have performed all our experiments on scenes "room_0", "room_1" and "office_0" since they represent a typical home and office environment. For the dataset set we have 2 different splitting options:
@@ 4.3 Architecture choice
There are several options of the architectures for our method which are shown in Fig.  2. To properly select the optimal hyperparameters $\lambda$ and $\omega$, we firstly have manually found rough estimates for the optimum, and then used 10 runs of Bayesian hyperparameter fitting for each architecture. We have chosen the search interval for optimal hyperparameters: $\sim Uniform[10^{-5},10^{-3}]$ for $\omega$ and $\sim Uniform[10^{-3},10^{-1}]$ for $\lambda$.
@@
Table 1: Hyperparameter values at which the best quality of reconstruction is achieved on a given architecture.
Network architecture
best $\lambda$
best $\omega$
mIoU
uSF (c0, s0)
[…]
uSF (c1, s0)
@@
[…]
0.5102
uSF (c0, s1)
[…]
0.5579
@@ 4.4 Positional encoding choice
We aimed to create an effective implementation of our method without deterioration in quality. To achieve this, we have used the hash positional encoding which is proposed in Instant-NGP [10] method. It allows us to significantly reduce the number of layers in the architecture and the size of dimensions. Hidden layer dimension is 64 instead of 256, number of dense layers is 5 instead of 8.
@@
Table 2: Influence of positional encoding and model size on the quality of reconstruction of 3D scenes from Replica dataset.
Scene
mIoU
PSNR
SSIM
LPIPS
Accuracy
$T_{render}$, s
Encoding
Architecture
room_0 0.961
31.024
0.902
0.109
0.980
1.993
hash
Small
0.749
22.538
0.667
0.562
0.823
0.893
freq
31.067 0.905 0.104
1.073
hash
22.486
0.666
0.567
1.011
freq
office_0 0.982
39.493
0.975
0.030
0.992
2.583
hash
Small
0.849
28.452
0.835
0.438
0.914
1.500
freq
39.958 0.979 0.022
1.557
hash
27.641
0.822
0.485
1.131
freq
room_1
36.369 0.956 0.042
0.753
hash
Small
33.179
0.896
0.164
10.300
freq
Large
@@ 4.5 Study of different uncertainties usage in uSF architecture
We have compared the recovery quality of neural radiance field and neural semantic field in cases where the model predicts only color and semantic labels, as well as when the model additionally estimates only rgb uncertainty, only semantic uncertainty and both types of uncertainty. The metric values are given in the Table 3. The experiments have been carried out on different train datasets sizes: small with 27 images and large with 562 images.
Table 3: Average values of mIoU and PSNR metrics for different for different numbers of images in the training dataset in cases where the model doesn’t quantify uncertainty, estimates only rgb uncertainty, only semantic uncertainty, and both types of uncertainties.
Training sample size Metrics
uSF (without
uncertainty)
uSF (rgb
uncertainty)
uSF (semantic
uncertainty)
uSF (both
uncertainty)
@@
As we can see, in the case where there are 27 images in the training set the average PSNR value increases slightly with the addition of the uncertainty estimation. The average mIoU value improves in the case of predicting only semantic uncertainty and both uncertainty. In the case of a large set, adding a prediction also has a positive effect on metric values. So adding an uncertainty estimation does not degrade the quality of the network’s predictions, and in some cases even improves it.
@@ 4.6 Influence of the $\omega$ hyperparameter on semantic field reconstruction
We tried to do a search and see how the quality of recovery depends on the parameter $\omega$. We fixed $\lambda=0.05$ and began to change $\omega$ and carry out experiments on two different scenes that capture ordinary room and office environment. The small set of training images was used. At first, we predicted only rgb uncertainty, after which we added prediction of semantic uncertainty. The results are shown in the Fig.  4.
Figure 4: The mean values of mIoU and PSNR depending on $\omega$ in the first row and $\lambda$ in the second row for selected scenes from Replica Dataset.
@@ 4.8 Quality of the semantic 3D-scene reconstruction
We have compared the neural semantic fields reconstruction quality of our uSF method trained with the large (562) set of images with the Semantic NeRF method on the scene room_1. Our method have showed the best metric values for both predicting semantic labels and predicting color.
Table 4: Quality of the semantic 3D-scene reconstruction of uSF and Semantic NeRF methods.
@@
Semantic NeRF  0.9313  0.2119 0.9526 0.9903

### v2

@@
Abstract. Recently, there has been an increased interest in NeRF methods which reconstruct differentiable representation of three-dimensional scenes. One of the main limitations of such methods is their inability to assess the confidence of the model in its predictions. In this paper, we propose a new neural network model for the formation of extended vector representations, called uSF, which allows the model to predict not only color and semantic label of each point, but also estimate the corresponding values of uncertainty. We show that with a small number of images available for training, a model that quantifies uncertainty performs better than a model without such functionality. Code of the uSF approach is publicly available at https://github.com/sevashasla/usf/.
Keywords: Neural semantic field, Neural network, Learning, Uncertainty, 3D scene
@@ 1 Introduction
Over the past few years, many researchers have focused on the development of the field of differentiable scene representation. This development began with the appearance of the first NeRF method [1], which demonstrated the photo-realistic quality of novel views generation. Various modifications of the original method allow us to additionally solve such problems as depth estimation [2], semantic segmentation[3, 4], object detection [5]. Among the disadvantages of the NeRF methods one can note high computational complexity [6, 7], the need for a large number of images [8, 9] in the training dataset and the inability to estimate the confidence of the model in its predictions. Quantifying uncertainty is a crucial aspect for minimizing possible risks for tasks in various fields, for example robotics [10] and autonomous driving [11, 12].
As shown in Fig. 1 we propose a novel NeRF-based method that reconstructs the neural radiance and the neural semantic fields. The previously proposed NeRF methods only considered estimating the uncertainty for color predictions, i.e. regression task [13, 14, 15]. Our method additionally introduces the possibility of quantifying uncertainty for predicted semantic labels, i.e. classification task. There are two types of uncertainty: aleatoric and epistemic. Aleatoric uncertainty reflects the noise inherent in the input data. Epistemic uncertainty reflects what the model does not know due to a lack of data, too simple architecture, or inefficient training process. In both cases, aleatoric uncertainty is estimated. The high value of uncertainty may also indicate errors occurred in the data preparation process, for example, poor quality of labelling in segmentation tasks. We choose aleatoric uncertainty over epistemic uncertainty because quantifying aleatoric uncertainty does not require major changes in network architecture and multiple inferences or training. Adding uncertainty estimation to the algorithm pipeline does not degrade the quality of reconstruction with a sufficient number of images and improves the quality with a small training dataset. Our method is also computationally efficient because it uses trainable positional encoding based on hashing [6]. We summarize our main contributions as follows:
  • We have proposed a new neural network model called uSF, which allows to add the ability to quantify the uncertainty for predicted neural semantic field, while maintaining quantifying uncertainty for neural radiance field.
@@
Fig. 1: We propose the model named uSF to predict both color and semantic labels and estimate corresponding uncertainty.
@@ 2 Related Work
2.1 Neural Radiance Fields
Due to the great popularity of the NeRF method, the researchers have developed a large number of works expanding the possibilities of the original method. Some papers address the inefficiency of training and rendering [6, 7, 16, 17]. The other methods make it possible to apply the NeRF under certain restrictions: limited number of images in the training dataset [8, 9, 18, 19], lack of information about camera poses [20], unlimited scene space [21], dynamic scenes [22, 23].
@@ 2.2 Neural Semantic Fields
There are NeRF methods that can predict not only the color at each point of the scene, but also corresponding semantic labels. The Semantic-NeRF [3] method reconstructs the neural semantic field with the help of an additional head in the model architecture. The process of getting semantic labels at a point is similar to getting a color. This method is highly robust to the presence of outliers in the training data, yielding quality results even with 90% noise.
Another approach to segmentation is shown in the NeRF-SOS [24] method. The work successfully combines self-supervised 2D visual features for 3D semantic segmentation. Separate head outputs for segmentation are trained similar to Dino-ViT [25]. NeSF [26] is the method that produces generalizable neural semantic fields. The authors propose to firstly recover 3D density fields upon which 3D semantic segmentation model is trained supervised by posed 2D semantic maps. The pre-trained NeRF model is used to sample the volumetric density grid for further conversion into a semantic-feature grid. Semantic Ray [4] also has generalization ability and uses Cross-Reprojection Attention module to obtain such result.
@@ 2.3 Uncertainty Estimation
To estimate the aleatoric uncertainty, a probability distribution is imposed on the outputs of the model. This approach assumes that there is noise in the input data, the distribution parameters of which can be predicted as some function of the data. To predict epistemic uncertainty, it is a common practice to impose a distribution on the model parameters (Bayesian neural networks (BNN) [27, 28, 29]). The disadvantage of these methods for quantifying epistemic uncertainty is the requirement to train the model several times and make multiple inferences. There are also approaches that simultaneously estimate both types of uncertainties [30, 31, 32].
Uncertainty quantification in various implementations has already been applied in several NeRF models. The Active NeRF  [8] is designed to deal with the limited number of images in the training set. In this method, the estimated aleatoric uncertainty is used for adding new frames to the training data during the iterations of active learning. In the NeRF-W  [13] method aleatoric uncertainty is used to identify non-static objects that should be removed from the prediction. This approach allows you to obtain the scene reconstruction consisting only of static objects. The two following methods focus on quantifying of epistemic uncertainty. In the S-NeRF  [14] method, it is assumed that all possible neural radiance fields for the scene are distributed according to a prior distribution, on the basis of which and the input data the model predicts a posterior distribution, with which the uncertainty is estimated. The CF-NeRF[15] method works similarly, but does not impose any preliminary restrictions on the distribution. Recently in Bayes’ Rays work [33] the authors have proposed a new method to estimate epistemic uncertainty that does not require any changes during training process. All of these methods try to quantify uncertainty for color or depth predictions, while in this paper we propose the approach for estimating corresponding uncertainty for semantic label predictions.
@@ 3 Methodology
Fig. 2: From left to right we show the predicted color, predicted semantic labels, rgb uncertainty and semantic uncertainty during training process.
@@ 3.1 Neural Network Model
Fig. 3: uSF architecture. Orange rectangles are Linear + ReLU layers. Yellow rectangles are outputs of our model. We study different variants of our architecture in terms of the color (c0-c3) and the semantic (s0-s1) branches.
The architecture of our model is MLP with ReLU activation functions and a hidden state dimension equal to 64 as shown in the Fig. 3. Similar to Active NeRF [8] and NeRF-W [13] pipeline we add an additional head to predict the color variance $\beta_{rgb}$ which we use as rgb uncertainty. In our work we assume that it depends on the view direction.
We add two heads for semantic segmentation task. One outputs mean values $(sem_{1},\dots,sem_{n})$, and the other predicts the variance $\beta_{sem}^{2}$. We consider outputs of semantic heads to depend only on coordinates of the point, but not on view direction.
@@ 3.3 Uncertainty for Neural Semantic Field
$\widehat{p_{i}}=\mathbb{E}\,\text{SoftMax}(\{logit_{1},\dots,logit_{n}\})_{i}$ (2)
@@
During calculation we use the "reparametrization trick". Let $\xi\sim\mathcal{N}(a,\sigma^{2})$, and $\eta\sim\mathcal{N}(0,1)$. Then $\xi\leavevmode\nobreak\ {\buildrel d\over{=}}\leavevmode\nobreak\ a+\sigma\cdot\eta$. Then to generate logits, we firstly generate random values from $\mathcal{N}(0,1)$, and then convert them to the form we want by multiplying by the standard deviation ($\beta_{sem}$) and adding the mean ($sem_{i}$). The example of the predicted semantic uncertainty is shown in the Fig. 2
@@ 3.4 Learning Approach
$L_{uncert}=\frac{1}{N}\sum_{i=1}^{N}\bigg{(}\frac{||C(r_{i})-rgb_{i}||^{2}}{2\beta_{rgb}(r_{i})^{2}}+\frac{1}{2}\log{\beta_{rgb}(r_{i})^{2}}+\frac{\eta}{N_{si}}\sum_{j=1}^{N_{si}}\alpha(r_{i}(t_{j}))\bigg{)},$ (7)
@@
$lr(i)=lr(0)\cdot 0.1^{\min\left(1,\frac{i}{N_{i}}\right)}\cdot\min\left(1,10^{-3}+\frac{i}{N_{wi}}\right),$ (8)
@@
We use an early stopping during training. We firstly calculate mean of last 3 LPIPS values on eval dataset and then if |current LPIPSmean(last 3 $|\text{current LPIPS}-\text{mean}(\text{last 3 LPIPS})|<10^{-3}$ is two times in a row, the training process is stopped.
@@ 4.1 Hardware and Software Setup
We have performed experiments on NVIDIA GeForce 2080 Ti, NVIDIA GeForce 2080, NVIDIA GeForce 3060 graphics cards. The model takes about 6GB of video memory during training. According to our estimates, full training lasts approximately 4-5 hours for one scene.
@@ 4.2 Datasets and Environments
We use a pre-processed dataset from the Semantic NeRF [3] method with image size of $640\times 480$. Each scene contains 900 images with different view directions. We have performed all our experiments on scenes "room_0 "room_1"and "office_0"since they represent a typical home and office environment. For the dataset we have 2 different splitting options:
@@ 4.3 Architecture Choice
There are several options of the architectures for our method which are shown in Fig. 3. The options differ in the different number of layers and the presence of dependence on the view direction. To properly select the optimal hyperparameters $\lambda$ and $\omega$, we firstly have manually found rough estimates for the optimum, and then used 10 runs of Bayesian hyperparameter fitting for each architecture. We have chosen the search interval for optimal hyperparameters: $\sim Uniform[10^{-5},10^{-3}]$ for $\omega$ and $\sim Uniform[10^{-3},10^{-1}]$ for $\lambda$.
@@
Network architecture best $\lambda$ best $\omega$ mIoU
@@
Tab. 1: Hyperparameter values at which the best quality of reconstruction is achieved on a given architecture.
@@ 4.4 Positional Encoding Choice
We aim to create an effective implementation of our method without deterioration in quality. To achieve this, we have used the hash positional encoding which is described in Instant-NGP [6] method. It allows us to significantly reduce the number of layers in the architecture and the size of dimensions. Hidden layer dimension is 64 instead of 256, number of dense layers is 5 instead of 8.
@@
Scene mIoU PSNR SSIM LPIPS Accuracy $\text{T}_{\text{render}}$, s Encoding Architecture
room_0 0.961 31.024 0.902 0.109 0.980 1.993 hash Small
0.749 22.538 0.667 0.562 0.823 0.893 freq
- 31.067 0.905 0.104 - 1.073 hash
- 22.486 0.666 0.567 - 1.011 freq
office_0 0.982 39.493 0.975 0.030 0.992 2.583 hash Small
0.849 28.452 0.835 0.438 0.914 1.500 freq
- 39.958 0.979 0.022 - 1.557 hash
- 27.641 0.822 0.485 - 1.131 freq
room_1 - 36.369 0.956 0.042 - 0.753 hash Small
- 33.179 0.896 0.164 - 10.300 freq Large
Tab. 2: Influence of positional encoding and model size on the quality of reconstruction of 3D scenes from Replica dataset.
@@ 4.5 Study of Different Uncertainties Usage in uSF Architecture
We have compared the recovery quality of neural radiance field and neural semantic field in cases where the model predicts only color and semantic labels, as well as when the model additionally estimates only RGB uncertainty, only semantic uncertainty and both types of uncertainty. The metric values are given in Table 3. The experiments have been carried out on different train datasets sizes: small with 27 images and large with 562 images.
Training sample size Metrics uSF (w/o uncert.) uSF (rgb uncert.) uSF (semantic uncert.) uSF (both uncert.)
@@
Tab. 3: Average values of mIoU and PSNR metrics for different numbers of images in the training dataset in cases where the model doesn’t quantify uncertainty, estimates only rgb uncertainty, only semantic uncertainty, and both types of uncertainties.
As we can see, in the case where there are 27 images in the training set the average PSNR value increases slightly with the addition of the uncertainty estimation. The average mIoU value improves in the case of predicting only semantic uncertainty and both uncertainties. In the case of a large set, adding a prediction also has a positive effect on metric values. So adding an uncertainty estimation does not degrade the quality of the network’s predictions, and in some cases even improves it.
Fig. 4: The mean values of mIoU and PSNR depending on $\omega$ in the first row and $\lambda$ in the second row for selected scenes from Replica Dataset.
@@ 4.6 Influence of the $\omega$ Hyperparameter on Semantic Field Reconstruction
We have performed experiments to see how the quality of recovery depends on the parameter $\omega$. We fix $\lambda=0.05$ and begin to change $\omega$ and carry out experiments on two different scenes that capture ordinary room and office environment. The small set of training images is used. At first, we predict only rgb uncertainty, after which we add prediction of semantic uncertainty.
The results are shown in the Fig. 4.
@@ 4.8 Quality of the Semantic 3D-Scene Reconstruction
We have compared the neural semantic fields reconstruction quality of our uSF method trained with the large (562) set of images with the Semantic NeRF method which is similar to our network without uncertainty estimations on the room_1 scene. Our method have showed the best metric values for both predicting semantic labels and predicting color.
@@
Tab. 4: Quality of the semantic 3D-scene reconstruction of uSF and Semantic NeRF methods.